# RefVTON: person-to-person Try on with Additional Unpaired Visual Reference

### v3

@@ 4.1 Quantative result
Table 2: Quantitative results on three subsets of the DressCode dataset [morelli2022dress]: upper body, lower body, and dresses. The best and second-best results are highlighted in bold and underline, respectively. The symbol “*” denotes results reported in prior work, while “+R” indicates results with reference image input, and “MF” refers to mask-free input images. The subscripts $p$ and $s$ denote specific evaluation metrics for precision and recall, respectively.
@@
Table 4 and Table 2 summarizes the quantitative evaluation on the DressCode dataset. Our method (RefVTON) outperforms all baselines, delivering higher try-on quality and achieving strong consistency with the target person’s pose and body structure. Integrating reference images (“+R”) further enhances the results, establishing a new state-of-the-art. Importantly, even in the mask-free setting—without agnostic masks or additional inputs—our method correctly preserves garment styles (e.g., clothing length and design) and maintains high pose alignment, while achieving accuracy comparable to or surpassing prior baselines, demonstrating robustness and practicality.
4.2 QUALITATIVE COMPARISON
Although the previous method can fit the garment onto a human body, the generated results often fail to accurately capture the garment’s cut, style, details, and other characteristics that determine how it should be worn. In contrast, our method achieves superior visual fidelity compared to baselines. As shown in Fig. 6, on the VITON dataset, our RefVTON produces a highly realistic rendering of challenging garment materials such as hollow or semi-transparent fabrics without relying on reference images. For example, in the case of garments with hollow lace structures, our model accurately preserves the delicate perforated patterns and material transparency. In contrast, the baselines fail to reconstruct these details, often generating incorrect solid textures or spurious dotted patterns. Moreover, our approach demonstrates the best preservation of garment patterns: the printed letters and logos on clothes remain clear and consistent with the input, highlighting its advantage in maintaining fine-grained appearance details. The generation quality is further enhanced with the addition of reference images, which intuitively verify the effectiveness of the proposed tryon method. Even without using agnostic masks and by directly transferring the clothing $\bm{c}_{i}$ onto person images $\bm{\bar{p}}_{i,\bm{c}_{j}}$, our method still outperforms most baseline approaches.
Figure 7: Qualitative comparison on the DressCode dataset., and the model is trained following the pipeline in Fig. 3 (b). “Ref.” denotes the additional reference $\bm{r}_{i}$ image is used during the inference, while “mask-free” indicates that the image is generated using an unmasked person $\bm{p}_{i}$ image instead of a masked agnostic image $\bm{a}_{i}$.
Our model also performs well on a benchmark comprising a wider variety of clothing types. As shown in Fig. 7, we evaluate our method on the DressCode dataset, where garments are categorized into three types: upper body, lower body, and dresses. Our approach produces more faithful and natural try-on results compared to previous methods. In particular, it renders reflective materials such as leather and metallic fabrics with superior realism, avoiding the over-smoothing or distortion artifacts commonly observed in other methods. Moreover, even without agnostic masks, our model can still perform consistent try-on guided by garment style, accurately preserving the length, structure, and overall design without introducing mismatched or inconsistent clothing shapes.
Table 3: We compare our approach with [xu2025ootdiffusion] under cross-dataset evaluation, reporting results on both paired and unpaired settings for the VITON-HD and DressCode datasets. The best results are highlighted in bold. The * marker refers to the results reported in previous work. The symbol “+R” denotes results with reference image input, and “MF” indicates mask-free input images. Subscripts $p$ and $u$ represent the paired and unpaired test settings, respectively.
Methods VITON-HD DressCode
[…]
EVTAR(Ours) 0.851 6.23 0.80 0.072 9.11 1.08 0.896 3.70 1.13 0.045 5.22 1.20
EVTAR+R(Ours) 0.859 5.13 0.62 0.060 8.59 0.87 0.903 3.14 0.97 0.038 5.03 1.11
EVTAR/MF(Ours) 8.88 0.82 5.03 1.23
EVTAR+R/MF(Ours) 8.39 0.65 4.87 1.10
4.3 Training on VRF Dataset and Evaluation
To better evaluate the effectiveness of our proposed person-to-person virtual try-on framework and reference data construction pipeline, we train our RefVTON model on the VRF dataset collected using our data engine introduced in sec 3.4. The training hyperparameters are consistent with those used in the DressCode and VITON-HD experiments. For evaluation, we quantitatively test the generated try-on images on the DressCode and VITON-HD test sets, with quantitative metrics reported in Table 3. Our method (RefVTON) consistently achieves superior or comparable performance to OOTDiffusion [xu2025ootdiffusion] across both VITON-HD and DressCode benchmarks, in both paired and unpaired settings. Although RefVTON is not explicitly trained on DressCode or VITON-HD, the model trained on the mixed VRF dataset outperforms dataset-specific baselines across most metrics (e.g., FID and LPIPS), demonstrating a strong generalization capability. These results highlight the effectiveness of our approach in learning transferable representations that maintain robustness across diverse datasets.
4.4 Ablation Study
5 Conclusion
This paper introduces RefTON, a virtual try-on framework that supports both mask-based and mask-free inference and leverages reference images to guide the try-on process. We extend Flux-Kontext to handle multi-condition inputs of varying resolutions via a modified positional index. To train RefTON, we propose a reference data generation pipeline integrating Qwen2.5-VL and Flux-Kontext. This design allows RefTON to faithfully preserve translucent fabrics, intricate designs, and fine details, consistently outperforming existing methods both quantitatively and qualitatively, achieving state-of-the-art performance in all settings.
Appendix A Appendix
A.1 Detailed Reference Image Generation Process
Firstly, the reference image should faithfully preserve the details of the target garment $\bm{c}_{i}$. This requires the editing model to have strong consistency abilities. The Flux-Kontext model has a robust ability to edit the target region corresponding to the text prompt while keeping unrelated regions—such as the area not related to the garment—unchanged. Moreover, the Flux-Kontext model can perform precise local editing according to the text prompt, in this case, focusing on the person under the garment. Therefore, we choose the popular Flux-Kontext model [labs2025flux1kontextflowmatching] to edit the input image conditioned on the text prompt. Specifically, we add a sentence such as “keep the {target cloth} cloth unchanged” in the text prompt.
Secondly, the the person in the reference image should look significantly different from the person in the target image. During training, we observed that if the reference image is too similar to the target image, the model tends to rely on a "shortcut"—directly copying from the reference image and ignoring the agnostic/person image $\bm{a}_{i}$$\bm{p}_{i}$ and the cloth $\bm{c}_{i}$. To avoid this, we ensure that the person in the reference image differs from the target image to better showcase the visual effect of the clothing when worn, rather than focusing on the appearance of the person themselves. To achieve this, we utilize the Text-to-Image (T2I) capabilities of the Flux-Kontext model. We extract an accurate description of the person’s appearance in the target image (e.g., "The model has an East Asian appearance, with light skin, long black hair, and a neutral expression…") and pass it as the negative prompt to the T2I model. In contrast, we provide an opposite description (e.g., "The model has an African appearance, with dark skin, short yellow hair, and a cheerful expression…") as the positive prompt to guide the editing of the target image, as shown in Figure 4(b).
Fig 9 illustrates selected text prompts from the prompt description bank used for reference image generation. Furthermore, Fig 10 exhibits a selection of the resulting reference data samples.
Figure 9: Text prompts from the Outfit and Action Description Bank. To ensure the model edits only the person while preserving the target clothing, we assign different outfits and action description categories to different clothing inputs.
Figure 10: Sample reference images generated by our reference data generation pipeline. The editing model takes the target person’s image as input and synthesizes corresponding reference images, while preserving the garment’s appearance to match the cloth.
A.2 Further Quantitative Evaluation on High Resolution
Here we provide more evaluation results of the generated image. To further evaluate our method under high-resolution settings, Table 4 reports quantitative results on both VITON-HD and DressCode at a resolution of 1024. Across the paired and unpaired protocols, RefTON and RefTON+R consistently achieve high performance on nearly all metrics. Notably, the mask-free variants (RefTON/MF and RefTON+R/MF) deliver particularly strong results. These results demonstrate that our framework scales effectively to high-resolution synthesis and remains robust across diverse virtual try-on settings.
@@ A.2 Further Quantitative Evaluation on High Resolution
Table 4: Quantitative comparison across VITON-HD [choi2021viton] and DressCode [morelli2022dress] at a resolution of 1024. The best and second best results are shown in bold and underline. “+R” denotes the use of reference images, and “MF” indicates mask-free inputs. Subscripts $p$ and $u$ represent the paired and unpaired test settings, respectively. Unless otherwise specified, the same notations carry the same meanings throughout the figures and tables in this paper.
@@
A.3 Additional Qualitative Results
In this section, we provide various visualizations to further demonstrate the robustness, generalization ability, and high-fidelity performance of our method across different datasets, clothing categories, and evaluation settings.
As shown in Fig. 11, 12, 13, 14, and 15, our approach consistently preserves garment details, structural correctness, and texture realism under both paired and unpaired scenarios, with or without mask-free (MF) inputs. Moreover, as illustrated in Fig.LABEL:fig:in-the-wild, even in challenging in-the-wild conditions, our model exhibits strong robustness—maintaining accurate body pose, preserving background integrity, and producing stable try-on results without introducing artifacts or unintended changes.
@@ A.3 Additional Qualitative Results
As shown in Fig. 11, our model faithfully reproduces complex and fine-grained clothing patterns. Even for garments with dense textures, irregular motifs, or high-frequency visual elements, the generated results retain clear, sharp, and recognizable patterns with minimal distortion. The strong pattern-preservation ability highlights the effectiveness of our approach in capturing both global appearance and subtle local details.
@@
Fig. 12 shows that our method handles garments with challenging structural designs, such as multi-layered regions, unique silhouettes, or uncommon shapes. The generated try-on results maintain correct garment geometry, coherent contours, and physically plausible spatial arrangements. This demonstrates that our framework models structural priors robustly, enabling accurate synthesis even under significant variations in shape.
@@
As shown in Figs. 13, 14, and 15, our approach performs consistently well across the three DressCode subsets. In the unpaired and mask-free settings, our model successfully preserves fabric materials, shading, and texture characteristics while achieving realistic alignment between the garment and human body. Across diverse clothing types—including tops, pants, skirts, and full-body dresses—the synthesized results maintain stable structure, smooth boundaries, and visually coherent integration, demonstrating strong generalization and robustness.
Figure 11: Qualitative paired results in VITON-HD dataset with complex patterns on clothes. “reference” denotes that a reference image is provided.
Figure 12: Qualitative paired results in VITON-HD dataset with complex structure on clothes. “reference” denotes that a reference image is provided.
Figure 13: Qualitative results of upper-body sub-set in Dresscode dataset unpaired setting. “reference” denotes that a reference image is provided, while “MF” indicates mask-free inputs using the original person image instead of a masked agnostic image.
Figure 14: Qualitative results of lower-body sub-set in Dresscode dataset unpaired setting. “reference” denotes that a reference image is provided, while “MF” indicates mask-free inputs using the original person image instead of a masked agnostic image.
Figure 15: Qualitative results of dresses sub-set in Dresscode dataset unpaired setting. “reference” denotes that a reference image is provided, while “MF” indicates mask-free inputs using the original person image instead of a masked agnostic image.

### v5

@@ Appendix B Further Quantitative Evaluation on High Resolution
Table S5: Quantitative results on three subsets of the DressCode dataset [morelli2022dress]: upper body, lower body, and dresses. The best and second-best results are highlighted in bold and underline, respectively. The symbol “*” denotes results reported in prior work, while “+R” indicates results with reference image input, and “MF” refers to mask-free input images. The subscripts $p$ and $s$ denote specific evaluation metrics for precision and recall, respectively.
@@
Table S6: Quantitative comparison across VITON-HD [choi2021viton] and DressCode [morelli2022dress] at a resolution of 1024. The best and second best results are shown in bold and underline. “+R” denotes the use of reference images, and “MF” indicates mask-free inputs. Subscripts $p$ and $u$ represent the paired and unpaired test settings, respectively. Unless otherwise specified, the same notations carry the same meanings throughout the figures and tables in this paper.
@@
To further evaluate our method under high-resolution settings, Table S6 reports quantitative results on both VITON-HD and DressCode at a resolution of 1024. Across the paired and unpaired protocols, RefTON and RefTON+R consistently achieve high performance on nearly all metrics. Notably, the mask-free variants (RefTON/MF and RefTON+R/MF) deliver particularly strong results. These results demonstrate that our framework scales effectively to high-resolution synthesis and remains robust across diverse virtual try-on settings.
Appendix C Additional Qualitative Results
In this section, we provide extensive qualitative visualizations to further demonstrate the robustness, generalization ability, and high-fidelity performance of our method across different datasets, clothing categories, and evaluation settings.
We present qualitative comparisons with methods that require additional text inputs, and explicitly annotate each method’s required inputs in Fig S11(a). Our method achieves the best visual quality with the fewest required inputs. Moreover, adding a reference image further improves intricate details (e.g., lace, transparency, and texture), highlighting the advantage of visual reference. Figure S11(b) shows qualitative results on the StreetTryOn dataset under various settings. Compared with the baselines, our model produces more faithful try-on details while better preserving the person’s pose and the background. We also provide qualitative comparisons under the mask-free setting, as shown in Fig. S11(c). Our model demonstrates the strongest mask-free virtual try-on capability among all compared methods.
Figure S11: (a)We gain better results with fewer inputs required on VITON. (b)StreetTryOn qualitative results (in-the-wild & P2P): our method yields better results.(c) Our model shows stronger mask-free capability.
As shown in Fig. S12, S13, S14, S15, and S16, our approach consistently preserves garment details, structural correctness, and texture realism under both paired and unpaired scenarios, with or without mask-free (MF) inputs. Moreover, as illustrated in Fig.1, even in challenging in-the-wild conditions, our model exhibits strong robustness—maintaining accurate body pose, preserving background integrity, and producing stable try-on results without introducing artifacts or unintended changes.
@@ Appendix C Additional Qualitative Results
As shown in Fig. S12, our model faithfully reproduces complex and fine-grained clothing patterns. Even for garments with dense textures, irregular motifs, or high-frequency visual elements, the generated results retain clear, sharp, and recognizable patterns with minimal distortion. The strong pattern-preservation ability highlights the effectiveness of our approach in capturing both global appearance and subtle local details.
@@
Fig. S13 further illustrates that our method handles garments with challenging structural designs, such as multi-layered regions, unique silhouettes, or uncommon shapes. The generated try-on results maintain correct garment geometry, coherent contours, and physically plausible spatial arrangements. This demonstrates that our framework models structural priors robustly, enabling accurate synthesis even under significant variations in shape.
@@
As shown in Figs. S14, S15, and S16, our approach performs consistently well across the three DressCode subsets. In the unpaired and mask-free settings, our model successfully preserves fabric materials, shading, and texture characteristics while achieving realistic alignment between the garment and human body. Across diverse clothing types—including tops, pants, skirts, and full-body dresses—the synthesized results maintain stable structure, smooth boundaries, and visually coherent integration, demonstrating strong generalization and robustness.
In-the-Wild results.
In addition to controlled benchmark evaluations, RefTON demonstrates strong robustness and generalization in challenging in-the-wild scenarios. As shown in Fig. 1, it produces high-quality try-on results under diverse poses, lighting conditions, and cluttered backgrounds. Our mask-free pipeline directly transfers garments without human parsing masks or pose estimators, while preserving body pose, global structure, and identity cues. Moreover, additional-reference try-on further improves garment geometry, texture details, and overall realism. RefTON also maintains strong background consistency and avoids unnecessary changes to non-garment regions, making it reliable for real-world applications.
Figure S12: Qualitative paired results in VITON-HD dataset with complex patterns on clothes. “reference” denotes that a reference image is provided.
Figure S13: Qualitative paired results in VITON-HD dataset with complex structure on clothes. “reference” denotes that a reference image is provided.
Figure S14: Qualitative results of upper-body sub-set in Dresscode dataset unpaired setting. “reference” denotes that a reference image is provided, while “MF” indicates mask-free inputs using the original person image instead of a masked agnostic image.
Figure S15: Qualitative results of lower-body sub-set in Dresscode dataset unpaired setting. “reference” denotes that a reference image is provided, while “MF” indicates mask-free inputs using the original person image instead of a masked agnostic image.
Figure S16: Qualitative results of dresses sub-set in Dresscode dataset unpaired setting. “reference” denotes that a reference image is provided, while “MF” indicates mask-free inputs using the original person image instead of a masked agnostic image.